OpenStudy (anonymous):

25% of a number is 75. A.18.75 B.300

OpenStudy (anonymous):

18.75

OpenStudy (michele_laino):

we can write; \[\frac{ 25 }{ 100 }N=75\] where N is your number, so: \[N=75*4=300\] so B.

OpenStudy (michele_laino):

more explanation: \[N=75*\frac{ 100 }{ 25 }=75*4=300\]

OpenStudy (michele_laino):

proof: \[\frac{ 25 }{ 100}*300=25*3=75\]

OpenStudy (jhannybean):

\[25\% =\frac{25}{100} = 0.25\]\[0.25x=75 \implies x = \frac{75}{0.25} = 300\]